The annual percentage yield (APY) for a savings account with a monthly compounding period is O not related to the nominal interest rate. O always lower than the nominal interest rate. 0 the same as the nominal interest rate. O always higher than the nominal interest rate.
Answer #1

Correct option is (4).

APY is also known as Effective interest rate (EAR), computed as

EAR = [1 + (r/N)]N - 1, where N: Number of compounding periods

For example, if nominal rate (r) = 15% = 0.15, then with monthly compounding, N = 12.

EAR = [1 + (0.15/12)]12 - 1 = (1 + 0.0125)12 - 1 = (1.0125)12 - 1 = 1.1608 - 1 = 0.1608 = 16.08% > 15%